Question 5.2a:
A

1350-kg

car is traveling with an initial speed of

26.0 m/s.



Determine the car's speed after

?2.15 ? 105 J

of net work is done on the car.
m/s

Question 5.2b:
A

11.6-kg

box is sliding across a horizontal floor. It has an initial speed of

1.40 m/s

and the only force acting on it is kinetic friction with magnitude

fk = 3.00 N.



Determine the distance the box will travel before coming to rest.
m
Answer #1

from work energy theorem

dKE = W


0.5*m*(vf^2 - vi^2) = -W

0.5*1350*(Vf^2 - (26*26)) = 2.15e5

Vf = 18.9 m/s <----answer

----------------


fk = -3

a = -fk/m = -3/11.6 = -0.258 m/s^2

vf^2 - vi^2 = 2*a*s

0 - 1.4*1.4 = -2*0.258*s


s = 3.79 m <------answer